OpenStudy (anonymous):

it takes emmet 25 minutes to finish 2/3 of his work. how long will it take to finish all of it

OpenStudy (anonymous):

I think that you divide 25 by 2 and then multiply the outcome by 3 to get your answer

OpenStudy (yuki):

exactly, all Work = (rate)*(time) problems work like that

OpenStudy (yuki):

the information tells us that 2/3 work = rate * 25min

OpenStudy (yuki):

so rate = 2/3 * 1/25 work/min

OpenStudy (yuki):

since you want to find time now, 1 work = (2/3*1/25 work/min) * time time = 1 * 3/2*25 min

OpenStudy (yuki):

so the idea is to find rate first, and then plug the rest of the numbers in the equation